# This Time is Different: An Observability Perspective on Time Series Foundation Models

### v1

@@ Abstract
We introduce Toto, a time series forecasting foundation model with 151 million parameters. Toto uses a modern decoder-only architecture coupled with architectural innovations designed to account for specific challenges found in multivariate observability time series data. Toto’s pre-training corpus is a mixture of observability data, open datasets, and synthetic data, and is 4-10$\times$ larger than those of leading time series foundation models. Additionally, we introduce Boom, a large-scale benchmark consisting of 350 million observations across 2,807 real-world time series. For both Toto and Boom, we source observability data exclusively from Datadog’s own telemetry and internal observability metrics. Extensive evaluations demonstrate that Toto achieves state-of-the-art performance on both Boom and on established general purpose time series forecasting benchmarks. Toto’s model weights, inference code, and evaluation scripts, as well as Boom’s data and evaluation code, are all available as open source under the Apache 2.0 License available at https://huggingface.co/Datadog/Toto-Open-Base-1.0 and https://github.com/DataDog/toto.
@@ 1 Introduction
Observability is the practice of collecting and analyzing data generated by distributed computer systems to detect, diagnose, and swiftly resolve performance and reliability issues [4]. A major component of observability is monitoring time series metrics; observability tools generate massive and diverse sets of metrics that reflect a system’s operational health over time. These metrics encompass a wide variety of indicators—such as memory usage, CPU load, disk I/O, network throughput, hit counts, error rates, and latency—that each exhibit distinct behavioral patterns, and collectively represent an important but under-studied subset of general time series data.
Accurately modeling observability metrics is essential for critical tasks like anomaly detection [5] (e.g., identifying spikes in error rates) and predictive forecasting [6] (e.g., anticipating resource exhaustion or scaling needs). Observability data present challenges for traditional forecasting methods due to diversity, high-dimensionality, and complex distributional characteristics (Section 4.3). Moreover, real-world observability systems routinely generate millions to billions of distinct time series [7, 8, 9], rendering fine-tuning or supervised training of complex models per time series infeasible. These operational challenges suggest a compelling use case for zero-shot time series foundation models (FMs). However, we find that existing FMs [10, 11, 12] trained for general-purpose forecasting struggle to generalize to observability data; see Fig. 1D and Section 5.
@@
    Toto (Time Series Optimized Transformer for Observability). We develop a novel open-weights time series forecasting foundation model, with a focus on zero-shot capabilities. Toto uses a modern decoder-only architecture coupled with architectural innovations to account for the specific challenges found in observability time series data: a novel per-variate patch-based causal scaling to address highly nonstationary sequences; proportional time-variate factorized attention to judiciously attend across a large number of covariates; and a Student-T mixture prediction head optimized via a robust composite loss to fit complex and highly skewed distributions. Toto’s pretraining corpus contains 4-10$\times$ more unique data points than those of other time series FMs [11, 12, 13, 14], using a mix of domain-specific observability time series data, multi-domain public datasets, and synthetic data.
@@
    Boom (Benchmark of Observability Metrics). We introduce an open-source benchmark specifically for observability time series. Boom includes a large-scale, novel dataset with 350 million observations across 2,807 distinct multivariate time series, approximately twice the size of the general-purpose GIFT-Eval benchmark [3]. Unlike existing multi-domain benchmarks, Boom is comprised entirely of real-world observability data. It captures diverse time series dynamics and challenging behaviors across several subdomains (Fig. 4), thus providing a uniquely realistic and comprehensive evaluation environment for forecasting models.
@@
Toto also achieves the top position by a wide margin on two standard general-purpose time series benchmarks—GIFT-Eval and Long Sequence Forecasting (LSF)—implying our observability-focused design also pays dividends in other time series domains [1, 15]. We additionally perform ablations to motivate Toto’s architecture design (Fig. 3B), and conduct data analyses of Boom to illustrate the unique aspects of observability data that pose significant modeling challenges (Section 4).
@@ 2 Related Work
Supervised models. Current observability systems typically rely on classical modeling approaches such as Holt-Winters [5], tree-based methods [16], or (S)ARIMA [17] for forecasting and anomaly detection [18]. These approaches require individual training for each dataset, impeding scalability in large real-world systems. While neural models have been shown to surpass these classical models in some settings [19] they are generally larger, more complex, and still require training for each datset; they are thus operationally infeasible in practice.
Time series foundation models. By pre-training on large multi-domain datasets, several time series foundation models [11, 12, 13, 14, 20, 21, 22, 23, 24, 25] have achieved impressive zero-shot prediction capabilities on general purpose time series benchmarks, eliminating the need for domain-specific training or finetuning. This approach is especially promising for observability workloads, as a single model can be deployed and horizontally scaled to provide low-latency and relatively low-cost zero-shot inference. Our evaluations indicate that Toto outperforms existing time series foundation models by a wide margin on both public forecasting benchmarks and on Boom (Section 5).
Time series benchmarks. Traditional benchmarks include Monash [26], LSF [2], M3 [27], and M4 [28]. These benchmarks are either commonly used for pretraining FMs (Monash) [11, 13] or are limited in capacity in measuring the impact of modern methods (LSF, M3, M4) [29]. Aksu et al. [1], Qiu et al. [30], and Ansari et al. [11] all recently introduced multi-domain benchmarks that are better suited in scale and complexity to evaluate general-purpose time series foundation models. GIFT-Eval [1] in particular is orders of magnitude larger than LSF (see Table 1 for details); introduces standardized evaluation protocols to facilitate fair comparisons across models; and is unique in containing a large decontaminated pretraining dataset for rigorous measurement of zero-shot capabilities without test data leakage.
@@
Challenges of observability data. Several works have explored observability time series data and the prospect of applying the foundation model paradigm to it. Joosen et al. [31] analyze serverless function logs from Huawei Cloud, highlighting challenges such as values spanning nine orders of magnitude, heavy-tailed distributions, and multi-scale seasonality. Toner et al. [32] find that existing time series foundation models perform poorly on a small dataset of cloud data. Woo et al. [33] pretrain models on CPU and memory metrics, while Palaskar et al. [34] propose an observability-specific architecture trained in a full-shot setting on four small-scale observability datasets. Both Boom and Toto’s pretraining corpus are more diverse and in most cases orders of magnitude larger than the datasets considered in prior studies. Toto is also the first observability-specialized foundation model that also performs competitively on broad benchmarks like GIFT-Eval and LSF.
@@ 3.1 Model architecture
Transformer models for time series forecasting have variously used encoder-decoder [2, 11, 15], encoder-only [13, 19, 35], and decoder-only architectures [12, 21]. Toto uses a decoder-only architecture (trained on a next-patch prediction task), as it has been shown to scale well with respect to training efficiency when provided with sufficient data [36, 37]. We use non-overlapping patch embeddings [19, 38, 39], with a patch of size $P$ = 64, to project input time series of context length $L$ = 4096 points to embeddings of size $64\times D$ per variate, where $D$ = 768 is the embedding dimension for our final model (Fig. 2B). We also utilize techniques demonstrated to yield performance and efficiency improvements in contemporary transformer literature, including pre-normalization [40], RMSNorm [41], SwiGLU feed-forward layers [42], and RoPE [43] with XPOS [44] for improved extrapolation.
@@
Patch-based causal instance normalization to handle highly nonstationary data. To improve generalization across varying input scales, instance normalization is commonly applied prior to embedding time series data (for example, RevIN [45]). However, computing normalization statistics from the entire series would leak information from future time steps. This violates the causality of next patch prediction training and results in poor performance (see ablation in Appendix E). Das et al. [12] normalize the entire series according to the statistics of the first patch. That approach preserves causality, but can be ineffective for highly nonstationary data with statistics that vary significantly over time, as is the case with observability data (see Section 4.3).
@@
where $x_{i}$ represents the input value and $w_{i}$ the corresponding weight at timestep $i$. We set the weight to 0 for padding positions and 1 for all other positions. We add a minimum value of 0.1 to the causal standard deviation, in order to limit the amount of scaling applied to any particular value and avoid numerical overflow. Timesteps within each patch share the normalization values determined by the final timestep of that patch. As computing causal statistics for every subsequence would have suboptimal $O(n^{2})$ complexity in the sequence dimension, we instead use Welford’s online algorithm [46], a method that provides numerically stable variance calculations in $O(n)$ time. We gain further efficiency with a vectorized adaptation of the algorithm, allowing for GPU parallelism.
@@
Proportional factorized attention to judiciously capture variate interactions. We design Toto to natively handle multivariate forecasting by analyzing relationships in the time dimension (“time-wise” interactions) and the variate dimension (“variate-wise” interactions). While prior works that do not utilize variate-wise relationships (such as PatchTST [19] and TimesFM [12]) can still achieve competitive performance on multivariate datasets, other studies (e.g. Woo et al. [13]) have shown benefits from including variate-wise attention in ablations. However, observability metrics are often high-cardinality, multivariate time series, and a full attention schema simultaneously attending to both the time and variate dimensions (as in Woo et al. [13]) can be computationally costly.
@@
Drawing from our experience that time relationships are often more important than cross-variate relationships, we propose a relaxation of factorized attention. Factorized attention strictly alternates attention operations in the time and variate dimensions, allowing for time and variate mixing with lower algorithmic complexity [47, 48, 49]. Our design provides more granular control over the relative proportion of time-wise and variate-wise interactions. Specifically, each transformer block has attention along only a single axis, and we can change the ratio of time-wise to variate-wise transformer blocks as a hyperparameter (as illustrated in Figure 2C). Toto uses an 11:1 ratio (11 time-wise transformer blocks followed by a single variate-wise transformer block), which we found via hyperparameter optimization (see Appendix A.6).
@@
Student-T mixture model (SMM) head to model heavy-tailed observability time series. Producing probabilistic outputs is a critical feature of time series models in several domains, including observability [50, 51, 52]. In order to produce probabilistic forecasts across the wide range of output distributions present in observability data, we employ a method based on Gaussian mixture models (GMMs), which can approximate any density function [53]. We found that fitting GMMs in the presence of the extreme outliers and high skew found in observability data (see Fig. 5) leads to numerical instability in training, so we instead utilize a Student-T mixture model (SMM) of $K$ distributions, which robustly generalizes GMMs [54] and has shown promise for modeling heavy-tailed financial time series [55, 56]. In a contemporaneous work, Yao et al. [57] also explored time series foundation models which model a Student-T mixture output. A complete mathematical formulation of our Student-T mixture model, including equations and parameterizations, is provided in Appendix A.3.
@@
Composite robust loss to stabilize training dynamics. Mixture models optimized via maximum likelihood are known to suffer from singularities [58] and cluster collapse [59]. We use a composite loss formulation that we find, in practice, mitigates these effects. During training, we optimize a next-patch prediction task, where the model’s objective is to predict the distribution of values in the next patch given all previous patches. Our training combines the standard negative log-likelihood loss, $\mathcal{L}_{\text{NLL}}$, and a general robust loss, $\mathcal{L}_{\text{Robust}(\alpha,\delta)}$ [60] (see Appendix A.5). The robust loss provides a unified framework that allows for smoothly interpolating between several common robust loss functions [61, 62, 63, 64, 65, 66, 67, 68, 69], using parameters $\alpha\in[-\infty,2]$ and $\delta>0$ (see Fig. 7). In our case, after hyperparameter optimization, we found the Cauchy loss ($\alpha=0$) performed best in our setting:
@@
$\mathcal{L}_{\text{Robust}(0,\delta)}(x_{t},\hat{x}_{t})=\mathcal{L}_{\text{Cauchy}}(x_{t},\hat{x}_{t},\delta)=\log\left(\frac{1}{2}((x_{t}-\hat{x}_{t})/\delta)^{2}+1\right)$
@@
with $\delta=0.1$. While the NLL loss utilizes the full probabilistic output of the model, the robust loss operates point-wise and measures the prediction error between the predicted SMM mean and the ground truth data point. The final combined loss used for training Toto is: $\mathcal{L}=\lambda_{\text{NLL}}\cdot\mathcal{L}_{\text{NLL}}+(1-\lambda_{\text{NLL}})\cdot\mathcal{L}_{\text{Robust}(\alpha,\delta)},$ where $\lambda_{\text{NLL}}\in[0,1]$ is a ratio tuned simultaneously with the robust loss hyperparameters, with optimal value $\lambda_{\text{NLL}}=0.57$. In summary, we find the composite loss critical for avoiding degenerate solutions and the robustness of the point-wise loss necessary for mitigating the effects of outliers that are common in observability data (see Section 4.3). Further details, including explicit definitions of each loss component, are provided in Appendix A.5.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Observability is the practice of collecting and analyzing data generated by distributed computer systems to detect, diagnose, and swiftly resolve performance and reliability issues [majors2022observability]. A major component of observability is monitoring time series metrics; observability tools generate massive and diverse sets of metrics that reflect a system’s operational health over time. These metrics encompass a wide variety of indicators—such as memory usage, CPU load, disk I/O, network throughput, hit counts, error rates, and latency—that each exhibit distinct behavioral patterns, and collectively represent an important but under-studied subset of general time series data.
Accurately modeling observability metrics is essential for critical tasks like anomaly detection [Ze_Gandalf] (e.g., identifying spikes in error rates) and predictive forecasting [chang_metric_forecasting] (e.g., anticipating resource exhaustion or scaling needs). Observability data present challenges for traditional forecasting methods due to diversity, high-dimensionality, and complex distributional characteristics (Section 4.3). Moreover, real-world observability systems routinely generate millions to billions of distinct time series [grafana_prometheus, cloudflare_prometheus, recruly_observability], rendering fine-tuning or supervised training of complex models per time series infeasible. These operational challenges suggest a compelling use case for zero-shot time series foundation models (FMs). However, we find that existing FMs [liu2024moirai, ansari2024chronoslearninglanguagetime, das2024a] trained for general-purpose forecasting struggle to generalize to observability data; see Fig. 1C and Section 5.
@@
[…]
@@
    Boom (Benchmark of Observability Metrics). We introduce an open-source benchmark specifically for observability time series. Boom includes a large-scale, novel dataset with 350 million observations across 2,807 distinct multivariate time series, approximately twice the size of the general-purpose GIFT-Eval benchmark [gifteval_leaderboard]. Unlike existing multi-domain benchmarks, Boom is comprised entirely of real-world observability data. It captures diverse time series dynamics and challenging behaviors across several subdomains (Fig. 4), thus providing a uniquely realistic and comprehensive evaluation environment for forecasting models.
@@
Toto also achieves the top position by a wide margin on two standard general-purpose time series benchmarks—GIFT-Eval and Long Sequence Forecasting (LSF)—implying our observability-focused design also pays dividends in other time series domains [aksu2024giftevalbenchmarkgeneraltime, Zhou2020]. We additionally perform ablations to motivate Toto’s architecture design (Fig. 3B), and conduct data analyses of Boom to illustrate the unique aspects of observability data that pose significant modeling challenges (Section 4).
@@ 2 Related Work
Supervised models. Current observability systems typically rely on classical modeling approaches such as Holt-Winters [Ze_Gandalf], tree-based methods [pmlr-v48-guha16], or (S)ARIMA [box1970time] for forecasting and anomaly detection [datadog_anomaly]. These approaches require individual training for each dataset, impeding scalability in large real-world systems. While neural models have been shown to surpass these classical models in some settings [Nie2023] they are generally larger, more complex, and still require training for each datset; they are thus operationally infeasible in practice.
Time series foundation models. By pre-training on large multi-domain datasets, several time series foundation models [ansari2024chronoslearninglanguagetime, das2024a, Woo2024, shi2025timemoe, garza2023timegpt1, rasul2023lagllama, gruver2023large, 10.5555/3692070.3693383, chen2024visionts, goswamiMoment] have achieved impressive zero-shot prediction capabilities on general purpose time series benchmarks, eliminating the need for domain-specific training or finetuning. This approach is especially promising for observability workloads, as a single model can be deployed and horizontally scaled to provide low-latency and relatively low-cost zero-shot inference. Our evaluations indicate that Toto outperforms existing time series foundation models by a wide margin on both public forecasting benchmarks and on Boom (Section 5).
Time series benchmarks. Traditional benchmarks include Monash [godahewa2021monash], LSF [Wu2021], M3 [MAKRIDAKIS2000451], and M4 [MAKRIDAKIS202054]. These benchmarks are either commonly used for pretraining FMs (Monash) [ansari2024chronoslearninglanguagetime, Woo2024] or are limited in capacity in measuring the impact of modern methods (LSF, M3, M4) [xu2024specialized]. aksu2024giftevalbenchmarkgeneraltime, qiu2024tfb, and ansari2024chronoslearninglanguagetime all recently introduced multi-domain benchmarks that are better suited in scale and complexity to evaluate general-purpose time series foundation models. GIFT-Eval [aksu2024giftevalbenchmarkgeneraltime] in particular is orders of magnitude larger than LSF (see Table 1 for details); introduces standardized evaluation protocols to facilitate fair comparisons across models; and is unique in containing a large decontaminated pretraining dataset for rigorous measurement of zero-shot capabilities without test data leakage.
@@
Challenges of observability data. Several works have explored observability time series data and the prospect of applying the foundation model paradigm to it. Joosen_2023 analyze serverless function logs from Huawei Cloud, highlighting challenges such as values spanning nine orders of magnitude, heavy-tailed distributions, and multi-scale seasonality. toner2025performancezeroshottimeseries find that existing time series foundation models perform poorly on a small dataset of cloud data. woo2023pushinglimitspretrainingtime pretrain models on CPU and memory metrics, while palaskar2023automixerimprovedmultivariatetimeseries propose an observability-specific architecture trained in a full-shot setting on four small-scale observability datasets. Both Boom and Toto’s pretraining corpus are more diverse and in most cases orders of magnitude larger than the datasets considered in prior studies. Toto is also the first observability-specialized foundation model that also performs competitively on broad benchmarks like GIFT-Eval and LSF.
@@ 3.1 Model architecture
Transformer models for time series forecasting have variously used encoder-decoder [Wu2021, ansari2024chronoslearninglanguagetime, Zhou2020], encoder-only [Woo2024, Nie2023, Liu2024], and decoder-only architectures [das2024a, rasul2023lagllama]. Toto uses a decoder-only architecture (trained on a next-patch prediction task), as it has been shown to scale well with respect to training efficiency when provided with sufficient data [Radford2018ImprovingLU, Radford2019LanguageMA]. We use non-overlapping patch embeddings [Nie2023, DBLP:conf/iclr/CordonnierLJ20, dosovitskiy2021an], with a patch of size $P=64$, to project input time series of context length $L=4096$ points to embeddings of size $64\times D$ per variate, where $D=768$ is the embedding dimension for our final model (Fig. 2B). We also utilize techniques demonstrated to yield performance and efficiency improvements in contemporary transformer literature, including pre-normalization [xiong2020on], RMSNorm [zhang-sennrich-neurips19], SwiGLU feed-forward layers [shazeer2020gluvariantsimprovetransformer], and RoPE [Su2024] with XPOS [sun2022a] for improved extrapolation.
@@
Patch-based causal instance normalization to handle highly nonstationary data. To improve generalization across varying input scales, instance normalization is commonly applied prior to embedding time series data (for example, RevIN [kim2022reversible]). However, computing normalization statistics from the entire series would leak information from future time steps. This violates the causality of next patch prediction training and results in poor performance (see ablation in Appendix E). das2024a normalize the entire series according to the statistics of the first patch. Such an approach preserves causality, but can be ineffective for highly nonstationary data with statistics that vary significantly over time, as is the case with observability data (see Section 4.3).
@@
where $x_{i}$ represents the input value and $w_{i}$ the corresponding weight at timestep $i$. We set the weight to 0 for padding positions and 1 for all other positions. We add a minimum value of 0.1 to the causal standard deviation, in order to limit the amount of scaling applied to any particular value and avoid numerical overflow. Timesteps within each patch share the normalization values determined by the final timestep of that patch. As computing causal statistics for every subsequence would have suboptimal $O(n^{2})$ complexity in the sequence dimension, we instead use Welford’s online algorithm [welford1962note], a method that provides numerically stable variance calculations in $O(n)$ time. We gain further efficiency with a vectorized adaptation of the algorithm, allowing for GPU parallelism.
@@
Proportional factorized attention to judiciously capture variate interactions. We design Toto to natively handle multivariate forecasting by analyzing relationships in the time dimension (“time-wise” interactions) and the variate dimension (“variate-wise” interactions). While prior works that do not utilize variate-wise relationships (such as PatchTST [Nie2023] and TimesFM [das2024a]) can still achieve competitive performance on multivariate datasets, other studies (e.g. Woo2024) have shown benefits from including variate-wise attention in ablations. However, observability metrics are often high-cardinality, multivariate time series, and a full attention schema simultaneously attending to both the time and variate dimensions (also in Woo2024) can be computationally costly.
@@
Drawing from our experience that time relationships are often more important than cross-variate relationships, we propose a relaxation of factorized attention. Factorized attention alternates attention operations in the time and variate dimensions, allowing for time and variate mixing with lower algorithmic complexity [zhang2023crossformer, Rao2021, Arnab2021, gao2024building]. While prior works typically apply both variate-wise and time-wise attention in each layer, our design provides more granular control over the relative proportion of time-wise and variate-wise interactions. Specifically, each transformer block has attention along only a single axis, and we can change the ratio of time-wise to variate-wise transformer blocks as a hyperparameter (as illustrated in Figure 2C). Toto uses an 11:1 ratio (11 time-wise transformer blocks followed by a single variate-wise transformer block), which we found via hyperparameter optimization (see Appendix A.6).
In Section D.4, we observe that Toto scales significantly better with increasing variates when compared to other TSFMs, consistently achieving the lowest CUDA time in its forward pass pass when the number variates is greater than 30.
@@
Student-T mixture model (SMM) head to model heavy-tailed observability time series. Producing probabilistic outputs is a critical feature of time series models in several domains, including observability [zhu2017deep, lee2023maat, hang2024robust]. In order to produce probabilistic forecasts across the wide range of output distributions present in observability data, we employ a method based on Gaussian mixture models (GMMs), which can approximate any density fitsction [Goodfellow-et-al-2016]. We found that fitting GMMs in the presence of the extreme outliers and high skew found in observability data (see Fig. 5) leads to numerical instability in training, so we instead utilize a SMM of $K$ distributions, which robustly generalizes GMMs [Peel2000], and has shown promise for modeling heavy-tailed financial time series [Meitz2018AMA, Wong2009]. The mixture distribution used by Woo2024 contains a Student-T component alongside other distributions, in contrast to our model which solely uses Student-T components. In a contemporaneous work, yao2025neuralscalinglawstime also explored time series foundation models with a Student-T mixture output. A complete mathematical formulation of our SMM, including equations and parameterizations, is provided in Appendix A.3.
@@
Composite robust loss to stabilize training dynamics. Mixture models optimized via maximum likelihood are known to suffer from singularities [bishop2006chapter9] and cluster collapse [eisner_gmm_optimization]. We use a composite loss formulation that we find, in practice, mitigates these effects. During training, we optimize a next-patch prediction task, where the model’s objective is to predict the distribution of values in the next patch given all previous patches. Our training combines the standard negative log-likelihood loss, $\mathcal{L}_{\text{NLL}}$, and a general robust loss, $\mathcal{L}_{\text{Robust}(\alpha,\delta)}$ [DBLP:journals/corr/Barron17] (see Appendix A.5). The robust loss provides a unified framework that allows for smoothly interpolating between several common robust loss functions [BLACK199675, Donald_Bayesian_1986, Aubert_Deterministic_1994, sun_secrets_2010, Dennis01011978, Leclerc1989, DBLP:conf/icip/CharbonnierBAB94, Huber1964RobustEO, ZHANG199759], using parameters $\alpha\in[-\infty,2]$ and $\delta>0$ (see Fig. 7). After hyperparameter optimization, we found the Cauchy loss ($\alpha=0$) performed best in our setting:
@@
$\mathcal{L}_{\text{Robust}(0,\delta)}(x_{t},\hat{x}_{t})=\mathcal{L}_{\text{Cauchy}}(x_{t},\hat{x}_{t},\delta)=\log\left(\frac{1}{2}((x_{t}-\hat{x}_{t})/\delta)^{2}+1\right)\$
@@
with $\delta=0.1$. While the NLL loss utilizes the full probabilistic output of the model, the robust loss operates point-wise and measures the prediction error between the predicted SMM mean and the ground truth data point. The final combined loss used for training Toto is: $\mathcal{L}=\lambda_{\text{NLL}}\cdot\mathcal{L}_{\text{NLL}}+(1-\lambda_{\text{NLL}})\cdot\mathcal{L}_{\text{Robust}},$ where $\lambda_{\text{NLL}}\in[0,1]$ is a ratio tuned simultaneously with the robust loss hyperparameters, with optimal value $\lambda_{\text{NLL}}=0.57$. In summary, we find the composite loss critical for avoiding degenerate solutions and the robustness of the point-wise loss necessary for mitigating the effects of outliers that are common in observability data (see Section 4.3). Further details, including explicit definitions of each loss component, are provided in Appendix A.5.